M Limited has an authorised share capital of R1 500 000 divided into 1 500 000 ordinary shares of R1 each. The issued share capital at 31 March 2017 was R500 000 which was fully paid, and had been issued at par. On 1 April 2017, the directors, in accordance with the company’s Articles, decided to increase the share capital of the company by offering a further 500 000 ordinary shares of R1 each at a price of R1.60 per share, payable as follows:
• On application, including the premium R0.85 per share
• On allotment R0.25 per share
• On first and final call on 3 August 2017 R0.50 per share

REQUIRED:
Write up the ledger accounts necessary to record these transactions in the books of M Limited.
1
Expert's answer
2019-04-29T10:34:29-0400

1.On application, including the premium R0.85 per share  

500000*0,85 = 425000

Bank Account Dr. 


Share Application Account Cr. 


Explanation of this Transaction :


Company gets liquid asset, so bank account will debit. Share application is the creditor account which we will transfer to share capital account allotment of shares to shareholders.


2.On allotment R0.25 per share

500000*0,25 = 125000

Share Application Account Dr.


Share Capital Account Cr.


Explanation of this Transaction :


Application money on allotted shares is transferred to share capital account. In simple words, we have transfer current liability into our fixed liability.


3.On first and final call on 3 August 2017 R0.50 per share

500000*0,5 = 250000

Bank Account Dr.


Share Call Account Cr.


Explanation of Transaction :


Company may get call money in 1st call, 2nd call and final calls. So, when and which call money will be due, we will transfer this call money to share capital account. When call money will receive, we will debit the bank account and credit the share call account.


http://www.svtuition.org/2013/01/accounting-for-share-capital.html